# On endomorphism algebras of silting complexes over hereditary abelian categories

### v2

@@ Abstract.
Let $\mathcal{E}$ be the class of finite-dimensional algebras isomorphic to endomorphism algebras of silting complexes over hereditary abelian categories. It is proved that the class $\mathcal{E}$ is closed under taking idempotent quotients, idempotent subalgebras and $\tau$-reduction. We also show that the proper class consisting of shod algebras is also closed under these operations. In addition, several classic classes of algebras—including laura, glued, weakly shod algebras—are proved to be closed under idempotent quotients, thereby generalizing a known result originally established for specific idempotents.
@@
silting reduction, $\tau$-reduction, quasi-silted algebra, idempentent quotient
@@ 1. Introduction
For a given algebra $A$, the study of its idempotent subalgebra $eAe$ and the idempotent quotient $A/AeA$ associated with an idempotent $e\in A$ is a central theme in representation theory (see, for instance, [CPS, AC, Xi, Xu, CK]). A primary motivation for focusing on these two constructions originates from the theory of recollements of derived categories. It is well known that under suitable conditions, the bounded derived category $\mathcal{D}^{b}(A)$ can be “glued” from $\mathcal{D}^{b}(eAe)$ and $\mathcal{D}^{b}(A/AeA)$ via a recollement [CPS]. This framework provides a powerful mechanism for homological reduction, allowing one to decompose the homological properties of AA—such as the finiteness of its global dimension—into those of $eAe$ and $A/AeA$ [AKLY].
@@
From this perspective, establishing the closure properties of a specific class of algebras under these operations is of fundamental importance. It transforms the class into a self-contained system where global conjectures and structural problems can be tackled through recursive arguments. For example, the fact that the class of $2$-Calabi-Yau tilted algebras arising from hereditary categories is closed under idempotent quotients was a key step in proving the connectedness of cluster-tilting graphs for hereditary categories [FG]. Similarly, the proof of the connectedness of $\tau$-tilting graphs for gentle algebras is based on the fact that the class of gentle algebras is closed under $\tau$-reduction [FGLZ].
@@
In the representation theory of finite-dimensional algebras, the study of endomorphism algebras of canonical objects—such as tilting modules, support $\tau$-tilting modules, and (2-term) silting complexes—is a central theme. These constructions define several fundamental classes of algebras, including tilted, quasi-tilted, silted, and quasi-silted algebras. A landmark result by Happel, Reiten, and Smalø [HRS] established that almost hereditary algebras are precisely the endomorphism algebras of tilting objects in hereditary abelian categories. This categorical characterization was further extended by Buan and Zhou [BZa, BZb], who demonstrated that shod algebras can be realized as endomorphism algebras of 2-term silting complexes over hereditary abelian categories. As a further generalization, endomorphism algebras of silting complexes over a finite-dimensional hereditary algebras were investigated in [ALLT] from the view point of structure of categories.
@@
The purpose of this work is to investigate the class $\mathcal{E}$ of finite-dimensional algebras realized as endomorphism algebras of basic silting complexes over hereditary abelian categories, specifically focusing on their behavior under the operations of taking idempotent subalgebras and idempotent quotients. Employing the machinery of silting reduction in triangulated categories, we develop an approach to analyze the stability of this class of algebras. Our main results, which are interconnected and proved using this method, establish the following closure properties.
@@
Let $A\in\mathcal{E}$ and $e\in A$ be an idempotent element. Then
@@
Note that the idempotent quotient can be viewed as a special case of the $\tau$-reduction introduced by Jasso [J]. It is therefore natural to consider whether the class $\mathcal{E}$ is closed under the more general operation of $\tau$-reduction.
@@
    The $\tau$-reduction of $A$ with respect to $Z$ belongs to $\mathcal{E}$.
@@
    If $A$ is quasi-silted, then the $\tau$-reduction of $A$ with respect to $Z$ is also quasi-silted.
The proof of Theorem 1.1, which establishes the closure property under idempotent quotients, naturally motivates a further investigation into the classical classes of algebras studied by Assem and Coelho [AC]. Specifically, we are led to consider whether these well-known classes—namely laura, glued, weakly shod, and shod algebras—are also closed under quotients by arbitrary idempotents. Utilizing an independent analysis with techniques distinct from silting reduction, we provide an affirmative answer to this question.
@@
Let $A$ be a finite-dimensional $k$-algebra and $e\in A$ be an idempotent element.
@@
The paper is organized as follows. We provide the necessary preliminaries and the main method of silting reduction in Section 2. In Section 3, we present the proof of Theorems 1.1-1.2. Section 4 is devoted to the proof of Theorem 1.3. In Section 5, we provide some examples clarifying the boundaries of our results.
@@
Throughout this paper, let $k$ be a field. By a hereditary abelian category, we mean a $\operatorname{\mathsf{Hom}}$-finite and $\operatorname{\mathsf{Ext}}$-finite $k$-linear hereditary abelian categry.
  • For a category $\mathcal{C}$ and a subcategory $\mathcal{X}$ of $\mathcal{C}$. We denote by $\mathcal{X}^{\perp}$ the full subcategory of $\mathcal{C}$ consisting of objects $X\in\mathcal{C}$ such that $\operatorname{\mathsf{Hom}}_{\mathcal{C}}(\mathcal{X},X)=0$. We define $^{\perp}\mathcal{X}$ dually.
  • We denote by $[1]$ the suspension functor of a triangulated category $\mathcal{T}$. For subcategories $\mathcal{X}$ and $\mathcal{Y}$ of $\mathcal{T}$, we denote by $\mathcal{X}*\mathcal{Y}$ the full subcategory consisting of objects $E\in\mathcal{T}$ such that there is a triangle $X\to E\to Y\to X[1]$ with $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$. Given an object $X$ we denote by $\operatorname{\mathsf{thick}}X$ the smallest thick subcategory of $\mathcal{T}$ containing $X$.
  • In a Krull-Schmidt category, for an object $X$, denote by $|X|$ the number of pairwise non-isomorphic indecomposable direct summands of $X$, and by $\operatorname{\mathsf{add}}\nolimits X$ the full subcategory of the finite direct sums of direct summands of $X$.
  • For an algebra $A$, we denote by $\operatorname{\mathsf{mod}}\nolimits A$ its category of finitely generated right modules, and by $\operatorname{\mathsf{ind}}A$ the set of isomorphism classes of indecomposable right $A$-modules. For an element $x\in A$, write $\langle x\rangle=AeA$.
@@ 2.1. Perpendicular subcategories in triangulated categories
Let $\mathcal{T}$ be a triangulated category and $\mathcal{X}$ a subcategory of $\mathcal{T}$. A morphism $f:X\to Y$ is called a right $\mathcal{X}$-approximation of $Y\in\mathcal{T}$, if $X\in\mathcal{X}$ and $\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},f):\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},X)\to\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},Y)$ is surjective. The subcategory $\mathcal{X}$ is contravariantly finite in $\mathcal{T}$, if each object in $\mathcal{T}$ has a right $\mathcal{X}$-approximation. Dually, we can define left $\mathcal{X}$-approximation and covariantly finite subcategories. The subcategory $\mathcal{X}$ is called functorially finite, if it is both contravariantly finite and covariantly finite.
@@
    If $\mathcal{S}$ is contravariantly finite, then the restriction $\mathcal{S}^{\perp}\to\mathcal{T}/\mathcal{S}$ of $\mathbb{L}$ is a triangle equivalence,
@@
    If $\mathcal{S}$ is covariantly finite, then the restriction $^{\perp}\mathcal{S}\to\mathcal{T}/\mathcal{S}$ of $\mathbb{L}$ is a triangle equivalence,
@@ 2.2. Silting theory
Let $\mathcal{T}$ be a $\operatorname{\mathsf{Hom}}$-finite Krull-Schmidt triangulated category. An object $T\in\mathcal{T}$ is called a presilting object if $\operatorname{\mathsf{Hom}}_{\mathcal{T}}(T,T[>0])=0$. A presilting object $T\in\mathcal{T}$ is silting if $\operatorname{\mathsf{thick}}T=\mathcal{T}$. The following result gives a characterization of a thick subcategory generated by a presilting object.
@@
There are triangulated categories which do not have silting objects. But if $\mathcal{T}$ admits a silting object, then the Grothendieck group $K_{0}(\mathcal{T})$ of $\mathcal{T}$ is a free abelian group of rank $|T|$ ([AI, Theorem 2.27]). In particular, each silting object of $\mathcal{T}$ has the same number of pairwise non-isomorphic indecomposable direct summands.
@@
where $f_{M}$ is a minimal left $\operatorname{\mathsf{add}}\nolimits\overline{T}$-approximation of $M$. According to [AI, Theorem 2.31], $N\oplus\overline{T}$ is a basic silting object of $\mathcal{T}$ and $\mu_{M}^{-}(T):=N\oplus\overline{T}$ is called the left mutation of $T$ with respect to $M$.
We denote by $\operatorname{\mathsf{silt}}\mathcal{T}$ the set of isomorphism classes of basic silting objects of $\mathcal{T}$.
@@
$\{T\in\operatorname{\mathsf{silt}}\ \mathcal{T}~|~D\in\operatorname{\mathsf{add}}\nolimits T\}\to\operatorname{\mathsf{silt}}\ \mathcal{T}/\mathcal{S}$
@@
Let us recall the inverse map of the bijection following the proof of [AI, Theorem 2.37]. Denote by $\mathcal{S}_{D}^{\leq 0}:=\cup_{l\geq 0}\operatorname{\mathsf{add}}\nolimits D*\operatorname{\mathsf{add}}\nolimits D[1]*\cdots*\operatorname{\mathsf{add}}\nolimits D[l]$ and $\mathcal{S}_{D}^{<0}:=\mathcal{S}_{D}^{\leq 0}[1]$. It is known that $\mathcal{S}_{D}^{<0}$ is covariantly finite in $\mathcal{T}$. Since $\mathcal{S}$ is functorially finite, we may identify $\mathcal{T}/\mathcal{S}$ with $\mathcal{S}^{\perp}$. Let $N\in\mathcal{S}^{\perp}$ be a silting object of $\mathcal{S}^{\perp}$. Consider the following triangle
@@
where $S_{N}[1]\in\mathcal{S}_{D}^{<0}$ and $g$ is a minimal left $\mathcal{S}_{D}^{<0}$-approximation. According to the proof of [AI, Theorem 2.37], $T_{N}\oplus D$ is a silting object of $\mathcal{T}$ such that $\mathbb{L}(T_{N}\oplus D)=\mathbb{L}(N)$.
@@ 2.3. Perpendicular subcategories in hereditary categories
Let $\mathcal{H}$ be a hereditary abelian category, that is, an abelian category satisfying $\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{2}(-,-)=0$. For any $E\in\mathcal{H}$, define
$E^{\perp_{[0,1]}}=\{X\in\mathcal{H}~|~\operatorname{\mathsf{Hom}}_{\mathcal{H}}(E,X)=0=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(E,X)\}.$
@@
$E^{\perp_{[0,1]}}$ is a hereditary abelian subcategory of $\mathcal{H}$. In particular $E^{\perp_{[0,1]}}$ is closed under images, kernels and cokernels.
@@
It is straightforward to check $E^{\perp_{[0,1]}}$ is closed under images, kernels and cokernels in $\mathcal{H}$, and hence it is an abelian subcategory of $\mathcal{H}$. We only need to show $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)=0$ for any $X,Y\in E^{\perp_{[0,1]}}$. We assume a 2-extension $\varepsilon$ in $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)$ which is represented by the following exact sequence in $E^{\perp_{[0,1]}}$
@@
Since $E^{\perp_{[0,1]}}$ is an abelian subcategory of $\mathcal{H}$, we obtain two exact sequences in $E^{\perp_{[0,1]}}$
@@
We assume that they represent $\varepsilon_{1}\in\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(\operatorname{\mathsf{im}}\nolimits f,Y)$ and $\varepsilon_{2}\in\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(X,\operatorname{\mathsf{im}}\nolimits f)$ respectively. Note that for $M,N\in E^{\perp_{[0,1]}}$, we have $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(M,N)=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(M,N)$. Applying $\operatorname{\mathsf{Hom}}_{E^{\perp_{[0,1]}}}(X,-)$ to (2.1) yields the following exact sequence
@@
By the definition of the connecting map $\eta$, we have $\varepsilon=\eta(\varepsilon_{2})$. On the other hand, since $\mathcal{H}$ is hereditary, $u_{*}$ is surjective, and hence $\eta=0$. It follows that $\varepsilon=0$. Since $\varepsilon$ is arbitrary, we have $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)=0$. ∎
Denote by ${\mathcal{D}}^{b}(\mathcal{H})$ the bounded derived category of $\mathcal{H}$. As a generalization of $E^{\perp_{[0,1]}}$ for $E\in\mathcal{H}$, we define the following subcategory for any $D\in{\mathcal{D}}^{b}(\mathcal{H})$
$D^{\perp_{\mathbb{Z}}}=\{M\in\mathcal{H}~|~\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D,M[\mathbb{Z}])=0\}.$
The following is a generalization of Lemma 2.6.
@@
$D^{\perp_{\mathbb{Z}}}$ is a hereditary abelian subcategory of $\mathcal{H}$. In particular, $D^{\perp_{\mathbb{Z}}}$ is closed under images, kernels and cokernels.
@@
Write $D=D_{1}\oplus D_{2}\oplus\cdots\oplus D_{r}$, where each $D_{i}$ is indecomposable. Then $D_{i}=X_{i}[n_{i}]$ for some indecomposable object $X_{i}\in\mathcal{H}$ and some $n_{i}\in\mathbb{Z}$. By definition,
@@
The assertion holds by Lemma 2.6. ∎
@@
Consider the inclusion $D^{\perp_{\mathbb{Z}}}\hookrightarrow\mathcal{H}$. By Proposition 2.7, this inclusion extends to a triangle functor $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to{\mathcal{D}}^{b}(\mathcal{H})$, which maps a complex to itself. Let $X,Y\in D^{\perp_{\mathbb{Z}}}$, by Proposition 2.7 again, we have
@@
It follows that the induced triangle functor $\iota$ is fully faithful. Thus, we may identify ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ as a full subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$.
Let $\mathcal{S}=\operatorname{\mathsf{thick}}{D}$. Clearly, a complex in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ is in $\mathcal{S}^{\perp}$. Conversely, given $X\in\mathcal{S}^{\perp}$, since $\mathcal{H}$ is hereditary, $X\cong\oplus_{i}H^{i}X[-i]\in{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$. In particular, the image of $\iota$ is $\mathcal{S}^{\perp}$. Hence, $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to(\operatorname{\mathsf{thick}}{D})^{\perp}$ is a triangle equivalence. ∎

### v3

@@ Abstract.
Let $\mathcal{E}$ be the class of finite-dimensional algebras isomorphic to endomorphism algebras of silting complexes over hereditary abelian categories. It is proved that the class $\mathcal{E}$ is closed under taking idempotent subalgebras, idempotent quotients and $\tau$-tilting reduction. We also show that the proper class consisting of quasi-silted algebras is also closed under these operations. In addition, several classical classes of algebras—including laura, glued, and weakly shod algebras—are proved to be closed under idempotent quotients, thereby generalizing a result originally established for specific idempotents.
@@
silting reduction, $\tau$-tilting reduction, quasi-silted algebra, idempontent quotient
@@ 1. Introduction
For a given algebra $A$, the study of its idempotent subalgebra $eAe$ and the idempotent quotient $A/AeA$ associated with an idempotent $e\in A$ is a central theme in representation theory (see, for instance, [CPS, AC, Xi, Xu, CK]). A primary motivation for focusing on these two constructions originates from the theory of recollements of derived categories. It is well known that under suitable conditions, the bounded derived category $\mathcal{D}^{b}(A)$ can be glued from $\mathcal{D}^{b}(eAe)$ and $\mathcal{D}^{b}(A/AeA)$ via a recollement [CPS]. This framework provides a powerful mechanism for homological reduction, allowing one to decompose the homological properties of AA—such as the finiteness of its global dimension—into those of $eAe$ and $A/AeA$ [AKLY].
@@
From this perspective, establishing the closure properties of a specific class of algebras under these operations is of fundamental importance. It transforms the class into a self-contained system where global conjectures and structural problems can be tackled through recursive arguments. For example, the fact that the class of $2$-Calabi-Yau tilted algebras arising from hereditary categories is closed under idempotent quotients was a key step in proving the connectedness of cluster-tilting graphs for hereditary categories [FG]. Similarly, the proof of the connectedness of $\tau$-tilting graphs for gentle algebras relies on the fact that the class of gentle algebras is closed under $\tau$-tilting reduction [FGLZ].
@@
In the representation theory of finite-dimensional algebras, the study of endomorphism algebras of distinguished objects—such as tilting modules, support $\tau$-tilting modules, and (2-term) silting complexes—lies at the heart of the field. These constructions give rise to several fundamental classes of algebras, including tilted, quasi-tilted, silted, and quasi-silted algebras. A landmark result by Happel, Reiten, and Smalø [HRS] established that quasi-tilted algebras are precisely the endomorphism algebras of tilting objects in hereditary abelian categories. This categorical characterization was further extended by Buan and Zhou [BZa, BZb], who proved that shod algebras can be realized as endomorphism algebras of 2-term silting complexes over hereditary abelian categories. As a further generalization, endomorphism algebras of silting complexes over finite-dimensional hereditary algebras were investigated in [ALLT] from the viewpoint of categorical structure.
@@
The purpose of this work is to investigate the class $\mathcal{E}$ of finite-dimensional algebras realized as endomorphism algebras of basic silting complexes over hereditary abelian categories, specifically focusing on their behavior under the operations of taking idempotent subalgebras and idempotent quotients. Employing the machinery of silting reduction in triangulated categories, we develop an approach to analyze the stability of this class of algebras. Our main results, which are interconnected and established via this method, verify the following closure properties.
@@
Let $A\in\mathcal{E}$ and $e\in A$ be an idempotent. Then
@@
Note that the idempotent quotient can be viewed as a special case of the $\tau$-tilting reduction introduced by Jasso [J]. It is therefore natural to ask whether the class $\mathcal{E}$ is closed under the more general operation of $\tau$-tilting reduction.
@@
    The $\tau$-tilting reduction of $A$ with respect to $Z$ belongs to $\mathcal{E}$.
@@
    If $A$ is quasi-silted, then its $\tau$-tilting reduction with respect to $Z$ is also quasi-silted.
The proof of Theorem 1.1, which establishes the closure property under idempotent quotients, motivates a further exploration into the classical classes of algebras studied by Assem and Coelho [AC]. Specifically, we examine whether these well-known classes—namely laura, glued, weakly shod, and shod algebras—are also closed under quotients by arbitrary idempotents. By employing an independent analysis with techniques distinct from silting reduction, we provide an affirmative answer to this question.
@@
Let $A$ be a finite-dimensional $k$-algebra and $e\in A$ be an idempotent.
@@
The paper is organized as follows. We provide the necessary preliminaries and the main method of silting reduction in Section 2. In Section 3, we present the proofs of Theorems 1.1 and 1.2. Section 4 is devoted to the proof of Theorem 1.3. Finally, Section 5 provides several examples to illustrate the boundaries of our results.
@@
Throughout this paper, let $k$ be a field. By a hereditary abelian category, we mean a $\operatorname{\mathsf{Hom}}$-finite and $\operatorname{\mathsf{Ext}}$-finite $k$-linear hereditary abelian category.
  • Let $\mathcal{X}$ be a subcategory of a category $\mathcal{C}$. We denote by $\mathcal{X}^{\perp}$ the full subcategory of $\mathcal{C}$ consisting of objects $X\in\mathcal{C}$ such that $\operatorname{\mathsf{Hom}}_{\mathcal{C}}(\mathcal{X},X)=0$. The subcategory $^{\perp}\mathcal{X}$ is defined dually.
  • We denote by $[1]$ the suspension functor of a triangulated category $\mathcal{T}$. For subcategories $\mathcal{X}$ and $\mathcal{Y}$ of $\mathcal{T}$, let $\mathcal{X}*\mathcal{Y}$ be the full subcategory consisting of objects $E\in\mathcal{T}$ such that there is a triangle $X\to E\to Y\to X[1]$ with $X\in\mathcal{X}$ and $Y\in\mathcal{Y}$. Given an object $X\in\mathcal{T}$, we denote by $\operatorname{\mathsf{thick}}X$ the smallest thick subcategory of $\mathcal{T}$ containing $X$.
  • In a Krull-Schmidt category, for an object $X$, denote by $|X|$ the number of pairwise non-isomorphic indecomposable direct summands of $X$, and by $\operatorname{\mathsf{add}}\nolimits X$ the full subcategory consisting of finite direct sums of direct summands of $X$.
  • For an algebra $A$, we denote by $\operatorname{\mathsf{mod}}\nolimits A$ the category of finitely generated right $A$-modules, and by $\operatorname{\mathsf{ind}}A$ the set of isomorphism classes of indecomposable right $A$-modules. For an element $x\in A$, let $\langle x\rangle=AxA$ be the two-sided ideal generated by $x$.
@@ 2.1. Perpendicular subcategories in triangulated categories
Let $\mathcal{T}$ be a triangulated category and $\mathcal{X}$ a subcategory of $\mathcal{T}$. A morphism $f:X\to Y$ is called a right $\mathcal{X}$-approximation of $Y\in\mathcal{T}$ if $X\in\mathcal{X}$ and the induced map $\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},f):\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},X)\to\operatorname{\mathsf{Hom}}_{\mathcal{T}}(\mathcal{X},Y)$ is surjective. The subcategory $\mathcal{X}$ is contravariantly finite in $\mathcal{T}$ if every object in $\mathcal{T}$ admits a right $\mathcal{X}$-approximation. Dually, one defines left $\mathcal{X}$-approximation and covariantly finite subcategories. The subcategory $\mathcal{X}$ is called functorially finite if it is both contravariantly finite and covariantly finite.
@@
    If $\mathcal{S}$ is contravariantly finite, then the restriction $\mathcal{S}^{\perp}\to\mathcal{T}/\mathcal{S}$ of $\mathbb{L}$ is a triangle equivalence.
@@
    If $\mathcal{S}$ is covariantly finite, then the restriction $^{\perp}\mathcal{S}\to\mathcal{T}/\mathcal{S}$ of $\mathbb{L}$ is a triangle equivalence.
@@ 2.2. Silting theory
Let $\mathcal{T}$ be a $\operatorname{\mathsf{Hom}}$-finite Krull-Schmidt triangulated category. An object $T\in\mathcal{T}$ is called presilting if $\operatorname{\mathsf{Hom}}_{\mathcal{T}}(T,T[>0])=0$. A presilting object $T\in\mathcal{T}$ is said to be silting if $\operatorname{\mathsf{thick}}T=\mathcal{T}$. A silting object $T$ is called tilting if it also satisfies $\operatorname{\mathsf{Hom}}_{\mathcal{T}}(T,T[<0])=0$. The following result provides a characterization of the thick subcategory generated by a presilting object.
@@
While there exist triangulated categories that do not possess silting objects, if $\mathcal{T}$ admits a silting object $T$, then its Grothendieck group $K_{0}(\mathcal{T})$ is a free abelian group of rank $|T|$ [AI, Theorem 2.27]. Consequently, all silting objects in $\mathcal{T}$ have the same number of pairwise non-isomorphic indecomposable direct summands.
@@
where $f_{M}$ is a minimal left $\operatorname{\mathsf{add}}\nolimits\overline{T}$-approximation of $M$. According to [AI, Theorem 2.31], the object $\mu_{M}^{-}(T):=N\oplus\overline{T}$ is again a basic silting object of $\mathcal{T}$, called the left mutation of $T$ with respect to $M$.
We denote by $\operatorname{\mathsf{silt}}\mathcal{T}$ the set of isomorphism classes of basic silting objects in $\mathcal{T}$.
@@
$\{T\in\operatorname{\mathsf{silt}}\ \mathcal{T}\mid D\in\operatorname{\mathsf{add}}\nolimits T\}\to\operatorname{\mathsf{silt}}\ \mathcal{T}/\mathcal{S}$
@@
Let us recall the inverse map of the aforementioned bijection, following the proof of [AI, Theorem 2.37]. Let $\mathcal{S}_{D}^{\leq 0}:=\cup_{l\geq 0}\operatorname{\mathsf{add}}\nolimits D*\operatorname{\mathsf{add}}\nolimits D[1]*\cdots*\operatorname{\mathsf{add}}\nolimits D[l]$ and $\mathcal{S}_{D}^{<0}:=\mathcal{S}_{D}^{\leq 0}[1]$. It is known that $\mathcal{S}_{D}^{<0}$ is covariantly finite in $\mathcal{T}$. Furthermore, as $\mathcal{S}$ is functorially finite, we can identify $\mathcal{T}/\mathcal{S}$ with $\mathcal{S}^{\perp}$. Let $N\in\mathcal{S}^{\perp}$ be a silting object in $\mathcal{S}^{\perp}$. Consider the following triangle:
@@
where $S_{N}[1]\in\mathcal{S}_{D}^{<0}$ and $g$ is a minimal left $\mathcal{S}_{D}^{<0}$-approximation of $N$. According to the proof of [AI, Theorem 2.37], $T_{N}\oplus D$ is a silting object in $\mathcal{T}$ such that $\mathbb{L}(T_{N}\oplus D)=\mathbb{L}(N)$.
@@ 2.3. Perpendicular subcategories in hereditary categories
Let $\mathcal{H}$ be a hereditary abelian category, i.e., an abelian category satisfying $\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{2}(-,-)=0$. For any object $E\in\mathcal{H}$, we define
$E^{\perp_{[0,1]}}=\{X\in\mathcal{H}\mid\operatorname{\mathsf{Hom}}_{\mathcal{H}}(E,X)=0=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(E,X)\}.$
@@
The subcategory $E^{\perp_{[0,1]}}$ is a hereditary abelian subcategory of $\mathcal{H}$. In particular $E^{\perp_{[0,1]}}$ is closed under taking images, kernels, and cokernels.
@@
One can readily verify that $E^{\perp_{[0,1]}}$ is closed under taking images, kernels, and cokernels in $\mathcal{H}$; consequently, it is an abelian subcategory of $\mathcal{H}$. It suffices to show that $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)=0$ for any $X,Y\in E^{\perp_{[0,1]}}$. Let $\varepsilon\in\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)$ be a $2$-extension represented by the following exact sequence in $E^{\perp_{[0,1]}}$:
@@
Since $E^{\perp_{[0,1]}}$ is an abelian subcategory of $\mathcal{H}$, we obtain the following two exact sequences in $E^{\perp_{[0,1]}}$:
@@
Let these sequences represent $\varepsilon_{1}\in\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(\operatorname{\mathsf{im}}\nolimits f,Y)$ and $\varepsilon_{2}\in\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(X,\operatorname{\mathsf{im}}\nolimits f)$, respectively. Recall that for any $M,N\in E^{\perp_{[0,1]}}$, $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{1}(M,N)=\operatorname{\mathsf{Ext}}_{\mathcal{H}}^{1}(M,N)$. Applying the functor $\operatorname{\mathsf{Hom}}_{E^{\perp_{[0,1]}}}(X,-)$ to (2.1) yields the following exact sequence
@@
By definition, $\varepsilon$ is the image of $\varepsilon_{2}$ under the connecting map $\eta$, i.e., $\varepsilon=\eta(\varepsilon_{2})$. On the other hand, since $\mathcal{H}$ is hereditary, the map $u_{*}$ is surjective, which implies $\eta=0$. It follows that $\varepsilon=0$. As $\varepsilon$ was chosen arbitrarily, we conclude that $\operatorname{\mathsf{Ext}}_{E^{\perp_{[0,1]}}}^{2}(X,Y)=0$. ∎
Let ${\mathcal{D}}^{b}(\mathcal{H})$ be the bounded derived category of $\mathcal{H}$. Generalizing the definition of $E^{\perp_{[0,1]}}$ to an arbitrary object $D\in\mathcal{D}^{b}(\mathcal{H})$, we consider the subcategory
$D^{\perp_{\mathbb{Z}}}=\{M\in\mathcal{H}\mid\operatorname{\mathsf{Hom}}_{{\mathcal{D}}^{b}(\mathcal{H})}(D,M[\mathbb{Z}])=0\}.$
The following result extends Lemma 2.6 to this setting.
@@
The subcategory $D^{\perp_{\mathbb{Z}}}$ is a hereditary abelian subcategory of $\mathcal{H}$. In particular, $D^{\perp_{\mathbb{Z}}}$ is closed under taking images, kernels, and cokernels.
@@
Let $D=D_{1}\oplus D_{2}\oplus\cdots\oplus D_{r}$ be the decomposition of $D$ into indecomposable objects. Then, each $D_{i}$ is of the form $X_{i}[n_{i}]$ for some indecomposable object $X_{i}\in\mathcal{H}$ and $n_{i}\in\mathbb{Z}$. By the definition of $D^{\perp_{\mathbb{Z}}}$, we have
@@
The assertion then follows immediately from Lemma 2.6. ∎
@@
Consider the inclusion $D^{\perp_{\mathbb{Z}}}\hookrightarrow\mathcal{H}$. In view of Proposition 2.7, this inclusion extends to a triangle functor $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to{\mathcal{D}}^{b}(\mathcal{H})$, which maps a complex to itself. For any objects $X,Y\in D^{\perp_{\mathbb{Z}}}$, it follows from Proposition 2.7 that
@@
This implies that the induced triangle functor $\iota$ is fully faithful. Thus, we may view ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ as a full subcategory of ${\mathcal{D}}^{b}(\mathcal{H})$.
Let $\mathcal{S}=\operatorname{\mathsf{thick}}{D}$. Clearly, any complex in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$ belongs to $\mathcal{S}^{\perp}$. Conversely, for any $X\in\mathcal{S}^{\perp}$, the fact that $\mathcal{H}$ is hereditary implies $X\cong\bigoplus_{i}H^{i}(X)[-i]$, which lies in ${\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})$. Consequently, the image of $\iota$ is exactly $\mathcal{S}^{\perp}$, and thus $\iota:{\mathcal{D}}^{b}(D^{\perp_{\mathbb{Z}}})\to(\operatorname{\mathsf{thick}}{D})^{\perp}$ is a triangle equivalence. ∎